OpenStudy (anonymous):

1.write y=-2/5x-3 in standard form using integers. a.2x+5y=-15 b.2x+y=-3 c.2x+5y=-3 d.x+y=-15 2.hieght | 67| |69| |70| |72| |74| |74| |78| |79| wieght |183| 201|206|220|226|240|253|255| what is the correlation coefficient of the set of data round your anwser to the nearest thousandth a. -0.946 b.0.596 c.0.035 d.0.981 3.temperature |67| |77| |83| |85| |89| |94| |96| |99| number of cones |403| 447|457|465|489|503|543|576| about how many ice cream cones would you expect the shop to sell if the temperature one day is 106 degrees? a.579 b.585 c.602 d.617

OpenStudy (anonymous):

3.correlation coefficient for your data set is 0.109 how confident can you be that your predicted value will be reasonably close to the actual value? a.i cant be confident at all; this is about as close to a random guess as you can get b.i can be a little confident; it might be close, or it might be way off c.i can be very confident; it will be close, but it probably wont be exact d.i can be certain that my predicted value will match the actual value exactly 4. Solve |y-1/4|=6 a.y =-25/4 or y= -23/4 b.y=-25/4 or y= 23/4 c. y= 25/4 or y= -23/4 d. y= 23/4 or y=-23/4